### v1

@@ 2 Notation and preliminaries
Given a matrix $A\in{\mathds{R}}^{l\times q}$, we use $A_{ik}$ to denote the entry at the $i$-th row and the $k$-th column of $A$, and use $A_{k}$ to denote the $k$-th column of $A$. The transpose of $A$ is denoted by $A^{\top}$. When $l=q$, we use $A^{-1}$ to denote the inverse of $A$ if it is nonsingular. Given a subset $S\subseteq\{1,\ldots,q\}$, we use $|S|$ to denote its cardinality and use $A_{S}$ to denote the sub-matrix of $A$ by eliminating all the columns that are not indexed by $S$ from $A$. For the given index sets $I\subseteq\{1,\ldots,l\}$ and $S\subseteq\{1,\ldots,q\}$, we use $A_{IS}$ to denote the $|I|\times|S|$ sub-matrix of $A$ by removing all the rows and columns not in $I$ and $S$, respectively. Given two matrices $A,B\in{\mathds{R}}^{l\times q}$, $A\circ B$ denotes their Hadamard product. The inner product of $A,B$ is defined by $\langle A,B\rangle={\sf trace}(AB)$, and $\|A\|=\sqrt{\langle A,A\rangle}$ is the Frobenius norm. We use $E$ to denote an all-ones matrix, whose dimension will be specified from the context. For a matrix $A\in{\mathcal{S}}^{p}$, we use $A\succ 0$ (or $A\succeq 0$) to say that $A$ is positive definite (or positive semidefinite). Moreover, $A\prec 0$ (or $A\preceq 0$) means that $-A\succ 0$ (or $-A\succeq 0$).
Given a set $C\subseteq{\mathcal{E}}$, we use ${\sf lin}(C)$ to denote the largest linear subspace contained in $C$ (the lineality space of $C$), ${\sf aff}(C)$ to denote the smallest linear subspace that contains $C$ (the affine hull of $C$), and ${\rm cl}(C)$ to denote the closure of $C$. The paratingent cone of $C$ at $\bar{z}$ is defined by
@@
$\widehat{\mathcal{N}}_{C}(\bar{z}):=\{v\in{\mathcal{E}}\mid\langle v,z-\bar{z}\rangle\leq o(\|z-\bar{z}\|)\ \forall\,z\in C\}$
@@
When $C$ is a closed convex set, one has $\widehat{\mathcal{N}}_{C}(\bar{z})$ coincides with ${\mathcal{N}}_{C}(\bar{z})$, and both of them are simply called the normal cone of $C$ at $\bar{z}$ (in convex analysis [37]), i.e.,
${\mathcal{N}}_{C}(\bar{z})=\begin{cases}\{v\in{\mathcal{E}}\mid\,\langle v,z-\bar{z}\rangle\leq 0\ \forall z\in{\mathcal{E}}\},&\mbox{if }\bar{z}\in C,\\ \emptyset,&\mbox{otherwise}.\end{cases}$
When $C$ is a closed convex set, the tangent cone ${\mathcal{T}}_{C}(\bar{z})$ of $C$ at $\bar{z}\in C$ can be defined by ${\mathcal{T}}_{C}(\bar{z}):=({\mathcal{N}}_{C}(\bar{z}))^{\circ}$ [40, Example 6.24].
Given a function $\psi:{\mathcal{E}}\to(-\infty,\infty]$ with $\bar{z}\in{\mathcal{E}}$ such that $\psi(\bar{z})$ is finite, according to [40, Theorem 8.9] the (limiting) subdifferential of $\psi$ at $\bar{z}$ can be defined as
@@
Given a set-valued mapping $\Psi:{\mathcal{E}}\rightrightarrows{\mathcal{F}}$ , we use ${\sf gph}\Psi\in{\mathcal{E}}\times{\mathcal{F}}$ to denote the the graph of $\Psi$, i.e.,
@@
$\Psi(z)\cap{\mathcal{V}}\subset\Psi(z^{\prime})+\kappa\|z-z^{\prime}\|\mathds{B}_{{\mathcal{F}}}\quad\forall\,z,z^{\prime}\in{\mathcal{U}},$
where $\mathds{B}_{{\mathcal{F}}}$ is the closed unit ball in ${\mathcal{F}}$. Moreover, it has been perfectly characterized by the Mordukhovich criterion [27] that, under the assumption that ${\sf gph}\Psi$ is locally closed around $(\bar{z};\bar{w})\in{\sf gph}\Psi$, the Aubin property of $\Psi$ holds at $\bar{z}$ for $\bar{w}$ if and only if ${\mathcal{D}}^{*}\Psi(\bar{z},\bar{w})(0)=\{0\}$.
@@ 2.1 Coderivative related to positive semidefinite cone
where $P\in{\mathds{R}}^{p\times p}$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix with non-increasing diagonal elements such that $\Lambda_{\alpha\alpha}\succ 0$ and $\Lambda_{\gamma\gamma}\prec 0$. Here $\alpha,\beta$ and $\gamma$ are three sets of indices with the cardinalities $|\alpha|$, $|\beta|$ and $|\gamma|$. One has $P_{\alpha}\in{\mathds{R}}^{p\times|\alpha|}$, $P_{\beta}\in{\mathds{R}}^{p\times|\beta|}$ and $P_{\gamma}\in{\mathds{R}}^{p\times|\gamma|}$. We use $\Pi_{{\mathcal{S}}_{+}^{p}}(A)$ to denote the metric projection of $A$ to the cone ${\mathcal{S}}_{+}^{p}$ (under the Frobenius norm). For convenience, we denote $A_{+}:=\Pi_{{\mathcal{S}}_{+}^{p}}(A)$ and $A_{-}:=A-A_{+}\in-{\mathcal{S}}^{p}_{+}$. Then, it is obvious that $|\alpha|$ and $|\gamma|$ are the ranks of $A_{+}$ and $A_{-}$, respectively. Recall that ${\mathcal{S}}_{+}^{p}$ is a closed convex cone, and one has
@@
Denote the set of all partitions of the index set $\beta$ by $\mathscr{P}_{\beta}$. Let ${\mathds{R}}_{\geq}^{|\beta|}$ be the set of all the vectors in ${\mathds{R}}^{|\beta|}$ whose components are arranged in non-increasing order, i.e.,
@@
$\Xi_{1}=\begin{pmatrix}E_{\beta_{+}\beta_{+}}&E_{\beta_{+}\beta_{0}}&(\Xi_{1})_{\beta_{+}\beta_{-}}\\ E^{\top}_{\beta_{+}\beta_{0}}&0&0\\ (\Xi_{1})^{\top}_{\beta_{+}\beta_{-}}&0&0\end{pmatrix},$ (2.6)
@@
[9, Theorem 3.1 &\&& Proposition 3.3] Suppose that $A\in{\mathcal{S}}^{p}$ has the eigenvalue decomposition in (2.4). Then, $U\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p}}(A_{+},A_{-})(V)$ if and only if
@@
$\Sigma_{ik}:=\frac{\max\{\lambda_{i},0\}-\max\{\lambda_{k},0\}}{\lambda_{i}-\lambda_{k}},\quad i,k\in\{1,\ldots,p\}.$
In addition, $\widetilde{U}_{\beta\beta}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta|}}(0,0)(\widetilde{V}_{\beta\beta})$ holds if and only if there exist a matrix $\Xi_{1}\in\Upsilon_{|\beta|}$ in (2.5) and an orthogonal matrix $O\in{\mathds{R}}^{|\beta|\times|\beta|}$ such that
@@
We make the following remark to Lemma 2.1.
@@
For the case that $A=0$ in Lemma 2.1, one has $|\beta|=p$. Thus, by taking $P$ as the identity matrix, one can get $\widetilde{V}_{\beta\beta}=V$. One can further take $\beta_{+}=\beta$ and $\beta_{-}=\beta_{0}=\emptyset$ to get a partition of $\beta$. In this case, one has from (2.6) and (2.7) that $\Xi_{1}=E_{\beta\beta}$ and $\Xi_{2}=0\in{\mathcal{S}}^{|\beta|}$. Then, one has that (2.10) holds with $\widetilde{U}_{\beta\beta}:=0\in{\mathcal{S}}^{|\beta|}$ for any orthogonal matrix $O\in{\mathds{R}}^{|\beta|\times|\beta|}$. Consequently, one has $0\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p}}(0,0)(V)$ for any $V\in{\mathcal{S}}^{p}$.
@@ 2.2 Second-order sufficient conditions
In this subsection, we briefly introduce the second-order (sufficient) optimality conditions of the NLSDP (1.8), which can be viewed as an extension of the discussions in [42].
Let $\bar{x}$ be a stationary point of the NLSDP (1.8) and $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\ldots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$ be a multiplier at $\bar{x}$, where $\bar{\zeta}\in{\mathds{R}}^{m}$ and $\bar{\Gamma}_{j}\in{\mathcal{S}}^{p_{j}}$ for all $j=1,\dots,J$. From the KKT system (1.3) we know that $g_{j}(\bar{x})\in{\mathcal{S}}_{+}^{p_{j}}$ and $\bar{\Gamma}_{j}\in{\mathcal{N}}_{{\mathcal{S}}_{+}^{p_{j}}}(g_{j}(\bar{x}))$. For convenience, define $g(\bar{x}):=(g_{1}(x);\ldots;g_{J}(\bar{x}))$ and $A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j}$, $j=1,\ldots,J$. Then one has $(A_{j})_{+}=g_{j}(\bar{x})$ and $(A_{j})_{-}=\bar{\Gamma}_{j}$. Moreover, one can write the eigenvalue decomposition of $A_{j}$ as in (2.4) with $P_{j}=((P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}})$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues. Then, the critical cone of (1.8) at $\bar{x}$ can be explicitly given as
${\mathcal{C}}(\bar{x})=\{d\,|\,h^{\prime}(\bar{x})d=0,\,g^{\prime}_{j}(\bar{x})d\in{\mathcal{C}}(A_{j};{\mathcal{S}}_{+}^{p_{j}}),\ j=1,\ldots,J\},$ (2.11)
where for any $j=1,\ldots,J$,
$\begin{array}[]{ll}{\mathcal{C}}(A_{j};{\mathcal{S}}_{+}^{p_{j}}):={\mathcal{T}}_{{\mathcal{S}}_{+}^{p_{j}}}\big{(}g_{j}(\bar{x})\big{)}\cap\bar{\Gamma}_{j}^{\perp}\\[5.69054pt] \ =\{Z\in{\mathcal{S}}^{p_{j}}\,|\ (P_{j})^{\top}_{\beta_{j}}Z(P_{j})_{\beta_{j}}\succeq 0,\,(P_{j})^{\top}_{\beta_{j}}Z(P_{j})_{\gamma_{j}}=0,\ (P_{j})^{\top}_{\gamma_{j}}Z(P_{j})_{\gamma_{j}}=0\}.\end{array}$ (2.12)
According to [6, Section 4.6.1], for a feasible point $x\in{\mathcal{X}}$ of (1.1), it is called nondegenerate [35], or the constraint nondegeneracy [36] holds at $x$, if
$G^{\prime}(x){\mathcal{X}}+{\sf lin}({\mathcal{T}}_{{\mathcal{K}}}(G(x)))={\mathcal{Y}},$ (2.13)
The constraint nondegeneracy condition (2.13) at $\bar{x}$ can be written as
@@
Note that (2.14) implies that ${\mathcal{M}}(\bar{x})$ is a singleton, i.e., $\bar{y}$ is the unique multiplier at $\bar{x}$. Meanwhile, by following the proof of [42, Proposition 3.1] one can get that
@@
As a consequence, the second-order sufficient condition and the SSOSC for the NLSDP problem (1.8) can be defined as follows.
@@
Let $\bar{x}$ be a stationary point of (1.8) with $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\dots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$. Let ${\mathcal{Q}}:{\mathcal{X}}\to{\mathcal{X}}$ be the self-adjoint linear operator defined by
$\langle d,{\mathcal{Q}}d\rangle=\langle d,\nabla_{xx}^{2}{\mathcal{L}}(\bar{x},\bar{y})d\rangle-2\sum_{j=1}^{J}\langle\bar{\Gamma}_{j},(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\rangle\quad\forall d\in{\mathcal{X}},$ (2.16)
where $(g_{j}(\bar{x}))^{\dagger}$ denotes the Moore–Penrose pseudoinverse of $g_{j}(\bar{x})$ for all $j=1,\ldots,J$. We say that the second-order sufficient condition holds at $(\bar{x};\bar{y})$ if
@@
In definition 2.1, the second-order sufficient condition (2.17) follows from [3], while the SSOSC is a straightforward extension of [42, Definition 3.2] from the case that $J=1$ to the general setting.

### v2

@@ 2 Notation and preliminaries
Given a matrix $A\in{\mathds{R}}^{l\times q}$, we use $A_{ik}$ to denote the entry at the $i$-th row and the $k$-th column of $A$, and use $A_{k}$ to denote the $k$-th column of $A$. The transpose of $A$ is denoted by $A^{\top}$. When $l=q$, we use $A^{-1}$ to denote the inverse of $A$ if it is nonsingular. Given a subset $S\subseteq\{1,\ldots,q\}$, we use $|S|$ to denote its cardinality and use $A_{S}$ to denote the sub-matrix of $A$ by eliminating all the columns that are not indexed by $S$ from $A$. For the given index sets $I\subseteq\{1,\ldots,l\}$ and $S\subseteq\{1,\ldots,q\}$, we use $A_{IS}$ to denote the $|I|\times|S|$ sub-matrix of $A$ by removing all the rows and columns not in $I$ and $S$, respectively. Given two matrices $A,B\in{\mathds{R}}^{l\times q}$, $A\circ B$ denotes their Hadamard product. The inner product of $A,B$ is defined by $\langle A,B\rangle={\sf trace}(A^{\top}B)$, and $\|A\|=\sqrt{\langle A,A\rangle}$ is the Frobenius norm. We use $E$ to denote an all-ones matrix, whose dimension will be specified from the context. For a matrix $A\in{\mathcal{S}}^{p}$, we use $A\succ 0$ (or $A\succeq 0$) to say that $A$ is positive definite (or positive semidefinite). Moreover, $A\prec 0$ (or $A\preceq 0$) means that $-A\succ 0$ (or $-A\succeq 0$).
Given a set $C\subseteq{\mathcal{E}}$, we use ${\sf lin}(C)$ to denote the largest linear subspace contained in $C$ (the lineality space of $C$), ${\sf aff}(C)$ to denote the smallest linear subspace that contains $C$ (the affine hull of $C$), ${\sf int}\,C$ to denote its topological interior, and ${\rm cl}\,C$ to denote the closure of $C$. The paratingent cone of $C$ at $\bar{z}$ is defined by
@@
$\widehat{\mathcal{N}}_{C}(\bar{z}):=\{v\in{\mathcal{E}}\mid\langle v,z-\bar{z}\rangle\leq o(\|z-\bar{z}\|)\ \forall\,z\in C\},$
@@
When $C$ is a closed convex set, $\widehat{\mathcal{N}}_{C}(\bar{z})$ coincides with ${\mathcal{N}}_{C}(\bar{z})$, and both of them are simply called the normal cone of $C$ at $\bar{z}$ (in convex analysis [40]), i.e.,
${\mathcal{N}}_{C}(\bar{z})=\begin{cases}\{v\in{\mathcal{E}}\mid\,\langle v,z-\bar{z}\rangle\leq 0\ \forall z\in C\},&\mbox{if }\bar{z}\in C,\\ \emptyset,&\mbox{otherwise}.\end{cases}$
When $C$ is a closed convex set, the tangent cone ${\mathcal{T}}_{C}(\bar{z})$ of $C$ at $\bar{z}\in C$ can be defined by ${\mathcal{T}}_{C}(\bar{z}):=({\mathcal{N}}_{C}(\bar{z}))^{\circ}$ (cf. [43, Example 6.24]).
Given a function $\psi:{\mathcal{E}}\to(-\infty,\infty]$ with $\bar{z}\in{\mathcal{E}}$ such that $\psi(\bar{z})$ is finite, according to [43, Theorem 8.9], the (limiting) subdifferential of $\psi$ at $\bar{z}$ is defined by
@@
Given a set-valued mapping $\Psi:{\mathcal{E}}\rightrightarrows{\mathcal{F}}$ , we use ${\sf gph}\Psi\subseteq{\mathcal{E}}\times{\mathcal{F}}$ to denote the graph of $\Psi$, i.e.,
@@
$\Psi(z)\cap{\mathcal{V}}\subseteq\Psi(z^{\prime})+\kappa\|z-z^{\prime}\|\mathds{B}_{{\mathcal{F}}}\quad\forall\,z,z^{\prime}\in{\mathcal{U}},$
where $\mathds{B}_{{\mathcal{F}}}$ is the closed unit ball in ${\mathcal{F}}$ centered at the origin. Moreover, such an Aubin property has been equivalently characterized by the Mordukhovich criterion [29]. Specifically, under the assumption that ${\sf gph}\Psi$ is locally closed around $(\bar{z};\bar{w})\in{\sf gph}\Psi$, the Aubin property of $\Psi$ holds at $\bar{z}$ for $\bar{w}$ if and only if ${\mathcal{D}}^{*}\Psi(\bar{z},\bar{w})(0)=\{0\}$.
@@ 2.1 Coderivative related to positive semidefinite cone
where $P\in{\mathds{R}}^{p\times p}$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix with non-increasing diagonal elements $\lambda_{1},\ldots,\lambda_{p}$ (the eigenvalues of $A$) such that $\Lambda_{\alpha\alpha}\succ 0$ and $\Lambda_{\gamma\gamma}\prec 0$. Here $\alpha,\beta$ and $\gamma$ are three sets of indices with the cardinalities $|\alpha|$, $|\beta|$ and $|\gamma|$. One has $P_{\alpha}\in{\mathds{R}}^{p\times|\alpha|}$, $P_{\beta}\in{\mathds{R}}^{p\times|\beta|}$ and $P_{\gamma}\in{\mathds{R}}^{p\times|\gamma|}$. We use $\Pi_{{\mathcal{S}}_{+}^{p}}(A)$ to denote the metric projection of $A$ to the cone ${\mathcal{S}}_{+}^{p}$ (under the Frobenius norm). For convenience, we denote $A_{+}:=\Pi_{{\mathcal{S}}_{+}^{p}}(A)$ and $A_{-}:=A-A_{+}\in-{\mathcal{S}}^{p}_{+}$. Then, it is obvious that $|\alpha|$ and $|\gamma|$ are the ranks of $A_{+}$ and $A_{-}$, respectively. Recall that ${\mathcal{S}}_{+}^{p}$ is a closed convex cone, and one has
@@
Note that $(A_{+},A_{-})\in{\sf gph}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p}}$, and in the following we will introduce the explicit formulation of the coderivative ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p}}(A_{+},A_{-})$. For this purpose, denote the set of all partitions of the index set $\beta$ by $\mathscr{P}_{\beta}$ and let ${\mathds{R}}_{\geq}^{|\beta|}$ be the set of all the vectors in ${\mathds{R}}^{|\beta|}$ whose components are arranged in non-increasing order, i.e.,
@@
$\Xi_{1}=\begin{pmatrix}E_{\beta_{+}\beta_{+}}&E_{\beta_{+}\beta_{0}}&(\Xi_{1})_{\beta_{+}\beta_{-}}\\[2.84526pt] E^{\top}_{\beta_{+}\beta_{0}}&0&0\\[2.84526pt] (\Xi_{1})^{\top}_{\beta_{+}\beta_{-}}&0&0\end{pmatrix},$ (2.6)
@@
[10, Theorem 3.1 &\&& Proposition 3.3] Suppose that $A\in{\mathcal{S}}^{p}$ has the eigenvalue decomposition in (2.4). Then, $U\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p}}(A_{+},A_{-})(V)$ if and only if
@@
$\Sigma_{ik}:=\frac{\max\{\lambda_{i},0\}-\max\{\lambda_{k},0\}}{\lambda_{i}-\lambda_{k}},\quad i,k\in\{1,\ldots,p\},$
where $/0$ 0 is defined to be $1$. In addition, $\widetilde{U}_{\beta\beta}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta|}}(0,0)(\widetilde{V}_{\beta\beta})$ holds if and only if there exist a matrix $\Xi_{1}\in\Upsilon_{|\beta|}$ in (2.5) and an orthogonal matrix $O\in{\mathds{R}}^{|\beta|\times|\beta|}$ such that
@@
We make the following remarks on Lemma 2.1.
@@
According to the definitions of $A_{+}$ and $A_{-}$, from (2.4) one can see that the index set $\beta$ corresponds to the failure of the strict complementarity. Specifically, the fact that $A_{+}\in{\mathcal{S}}^{p}_{+}$, $A_{-}\in{\mathcal{S}}^{p}_{-}$ and $\langle A_{+},A_{-}\rangle=0$ constitutes a complementarity condition, and this complementarity is strict if $A$ is not singular. When $|\beta|\neq 0$, an extensive non-smooth analysis should be imposed to establish the relation between $\widetilde{U}_{\beta\beta}$ and $\widetilde{V}_{\beta\beta}$ in (2.8), resulting in a more complicated coderivative inclusion (2.9) than the relationship for the other matrix pairs.
Remark 2.2.
For the case that $A=0$ in Lemma 2.1, one has $|\beta|=p$. Thus, by taking $P$ as the identity matrix, one can get $\widetilde{V}_{\beta\beta}=V$. In addition, one can further take $\beta_{+}=\beta$ and $\beta_{-}=\beta_{0}=\emptyset$ to get a partition of $\beta$. In this case, from (2.6) and (2.7) one has $\Xi_{1}=E_{\beta\beta}$ and $\Xi_{2}=0\in{\mathcal{S}}^{|\beta|}$. Then, one has that (2.10) holds with $\widetilde{U}_{\beta\beta}:=0\in{\mathcal{S}}^{|\beta|}$ for any orthogonal matrix $O\in{\mathds{R}}^{|\beta|\times|\beta|}$. Consequently, one has $0\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p}}(0,0)(V)$ for any $V\in{\mathcal{S}}^{p}$.
Furthermore, to understand Lemma 2.1 more intuitively, one can consider the following example, which is constructed from [10, Example 7.1].
Example 2.1.
Consider the matrix
$A=\begin{pmatrix}0&-2&-1\\ -2&0&-1\\ -1&-1&-1\end{pmatrix}\text{ with }A_{+}=\begin{pmatrix}1&-1&0\\ -1&1&0\\ 0&0&0\end{pmatrix}\text{ and }A_{-}=\begin{pmatrix}-1&-1&-1\\ -1&-1&-1\\ -1&-1&-1\end{pmatrix}.$
We have
$A=P\begin{pmatrix}2&0&0\\ 0&0&0\\ 0&0&-3\end{pmatrix}P^{\top}\text{ with }P=\begin{pmatrix}\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}&\frac{1}{\sqrt{3}}\\ -\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}&\frac{1}{\sqrt{3}}\\ 0&-\frac{2}{\sqrt{6}}&\frac{1}{\sqrt{3}}\end{pmatrix},$
and that the index sets of positive, zero, and negative eigenvalues are given by $\alpha=\{1\}$, $\beta=\{2\}$, and $\gamma=\{3\}$. Also, we have $\Sigma_{13}=\frac{2}{5}$. Since $|\beta|=1$ and $O\in\{1,-1\}$, three cases should be considered:
$\beta=\beta_{+}$, $\Xi_{1}=1$, and $\Xi_{2}=0$. One has $\widetilde{U}_{22}=0$.
$\beta=\beta_{0}$ and $\Xi_{1}=\Xi_{2}=0$. From (2.10) one has $\widetilde{U}_{22}\leq 0$ and $\widetilde{V}_{22}\leq 0$.
$\beta=\beta_{-}$, $\Xi_{1}=0$, and $\Xi_{2}=1$. One has $\widetilde{V}_{22}=0$.
Therefore, $U\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{3}}(A_{+},A_{-})(V)$ if and only if
$V=P\begin{pmatrix}\widetilde{V}_{11}&\widetilde{V}_{12}&\widetilde{V}_{13}\\ \widetilde{V}_{21}&\widetilde{V}_{22}&0\\ \widetilde{V}_{31}&0&0\end{pmatrix}P^{\top}\quad\mbox{and}\quad U=P\begin{pmatrix}0&0&\widetilde{U}_{13}\\ 0&\widetilde{U}_{22}&\widetilde{U}_{23}\\ \widetilde{U}_{31}&\widetilde{U}_{32}&\widetilde{U}_{\gamma\gamma}\end{pmatrix}P^{\top}$
with $2\widetilde{U}_{13}=3\widetilde{V}_{13}$ and $(\widetilde{U}_{22};\widetilde{V}_{22})\in({\mathds{R}}\times\{0\})\cup(\{0\}\times{\mathds{R}})\cup({\mathds{R}}_{-}\times{\mathds{R}}_{-})$.
@@ 2.2 Second-order sufficient conditions
In this subsection, we briefly introduce the second-order (sufficient) optimality conditions of the NLSDP problem (1.8), which can be viewed as an extension of the discussions in [45].
Let $\bar{x}$ be a stationary point of the NLSDP (1.8) and $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\ldots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$ be a multiplier at $\bar{x}$, where $\bar{\zeta}\in{\mathds{R}}^{m}$ and $\bar{\Gamma}_{j}\in{\mathcal{S}}^{p_{j}}$ for all $j=1,\dots,J$. From the KKT system (1.3) we know that $g_{j}(\bar{x})\in{\mathcal{S}}_{+}^{p_{j}}$ and $\bar{\Gamma}_{j}\in{\mathcal{N}}_{{\mathcal{S}}_{+}^{p_{j}}}(g_{j}(\bar{x}))$. For convenience, define
$g(\bar{x}):=(g_{1}(x);\ldots;g_{J}(\bar{x}))\quad\mbox{and}\quad A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j},\quad j=1,\ldots,J.$
Then one has $(A_{j})_{+}=g_{j}(\bar{x})$ and $(A_{j})_{-}=\bar{\Gamma}_{j}$. One also has the eigenvalue decomposition $A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ as in (2.4) with $P_{j}=((P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}})$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues.
Recall that at a stationary point $x\in{\mathcal{X}}$ of (1.1) with ${\mathcal{M}}(x)$ being nonempty, the critical cone at $x$ is defined by
${\mathcal{C}}(x):=\{d\in{\mathcal{X}}\mid G^{\prime}(x)d\in{\mathcal{T}}_{{\mathcal{K}}}(G(x)),\,f^{\prime}(x)d=0\}.$
Then, following the discussions in [45, Section 3], the critical cone of (1.8) at $\bar{x}$ can be explicitly given as
${\mathcal{C}}(\bar{x})=\big{\{}d\,|\,h^{\prime}(\bar{x})d=0,\,g^{\prime}_{j}(\bar{x})d\in{\mathcal{C}}(A_{j};{\mathcal{S}}_{+}^{p_{j}}),\ j=1,\ldots,J\big{\}},$ (2.11)
where for each $j=1,\ldots,J$,
$\begin{array}[]{ll}{\mathcal{C}}(A_{j};{\mathcal{S}}_{+}^{p_{j}}):={\mathcal{T}}_{{\mathcal{S}}_{+}^{p_{j}}}\big{(}g_{j}(\bar{x})\big{)}\cap\bar{\Gamma}_{j}^{\perp}\\[5.69054pt] \ =\big{\{}Z\in{\mathcal{S}}^{p_{j}}\,|\ (P_{j})^{\top}_{\beta_{j}}Z(P_{j})_{\beta_{j}}\succeq 0,\,(P_{j})^{\top}_{\beta_{j}}Z(P_{j})_{\gamma_{j}}=0,\ (P_{j})^{\top}_{\gamma_{j}}Z(P_{j})_{\gamma_{j}}=0\big{\}}.\end{array}$ (2.12)
According to [7, Section 4.6.1], for a feasible point $x\in{\mathcal{X}}$ of (1.1), it is called nondegenerate [38], or the constraint nondegeneracy [39] is called to hold at $x$, if
$G^{\prime}(x){\mathcal{X}}+{\sf lin}({\mathcal{T}}_{{\mathcal{K}}}(G(x)))={\mathcal{Y}}.$ (2.13)
Then for the NLSDP (1.8), the constraint nondegeneracy condition (2.13) at $\bar{x}$ can be written as
@@
Note that (2.14) implies that ${\mathcal{M}}(\bar{x})$ is a singleton, i.e., $\bar{y}$ is the unique multiplier at $\bar{x}$. In this case, by following the proof of [45, Proposition 3.1] one can get from (2.11) that
@@
Next, we discuss the second-order optimality conditions for (1.8). With $\bar{x}$ being a stationary point of (1.8) and $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\dots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$, we define the self-adjoint linear operator ${\mathcal{Q}}:{\mathcal{X}}\to{\mathcal{X}}$ by
$\langle d,{\mathcal{Q}}d\rangle=\langle d,\nabla_{xx}^{2}{\mathcal{L}}(\bar{x},\bar{y})d\rangle-2\sum_{j=1}^{J}\langle\bar{\Gamma}_{j},(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\rangle,\quad d\in{\mathcal{X}},$ (2.16)
where $(g_{j}(\bar{x}))^{\dagger}$ denotes the Moore–Penrose pseudoinverse of $g_{j}(\bar{x})$ for all $j=1,\ldots,J$. Based on (2.16), the second-order sufficient condition and the strong second-order sufficient condition (SSOSC) for (1.8) are defined as follows.
@@
Let $\bar{x}$ be a stationary point of (1.8) with $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\dots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$ and ${\mathcal{C}}(\bar{x})$ be the critical cone at $\bar{x}$. Let ${\mathcal{Q}}:{\mathcal{X}}\to{\mathcal{X}}$ be the self-adjoint linear operator defined by (2.16). We say that the second-order sufficient condition holds at $(\bar{x};\bar{y})$ if
@@
In Definition 2.1, the second-order sufficient condition given by (2.17) comes from [4], while the SSOSC is a straightforward extension and simplification of [45, Definition 3.2] from the case that $J=1$ to the general setting. It is the same as [45, Definition 3.2] when $J=1$ and the constraint nondegeneracy holds.
The following result on the relationship between the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ and the second-order sufficient condition (2.17) comes from [11, Corollary 25]. It constitutes the starting point of our analysis.
Lemma 2.2.
Let $\bar{x}$ be a locally optimal solution to (1.8) with $\bar{y}\in{\mathcal{M}}(\bar{x})$. Suppose that the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) holds at $(0;0)$ for $(\bar{x};\bar{y})$. Then, $\bar{x}$ is nondegenerate, in the sense that (2.14) holds, and the second-order sufficient condition (2.17) in Definition 2.1 holds with ${\mathcal{Q}}$ being defined in (2.16).